# A Mechanistic Interpretation of Syllogistic Reasoning in Auto-Regressive Language Models

### v1

@@ Abstract
Recent studies on logical reasoning in auto-regressive Language Models (LMs) have sparked a debate on whether such models can learn systematic reasoning principles during pre-training or merely exploit superficial patterns in the training data. This paper presents a mechanistic interpretation of syllogistic reasoning in LMs to further enhance our understanding of internal dynamics. Specifically, we present a methodology for circuit discovery aimed at disentangling content-independent reasoning mechanisms from world knowledge acquired during pre-training. Through two distinct intervention methods, we uncover a sufficient and necessary circuit involving middle-term suppression that elucidates how LMs transfer information to derive valid conclusions from premises. Furthermore, we investigate how belief biases manifest in syllogistic reasoning, finding evidence of partial contamination from additional attention heads responsible for encoding commonsense and contextualized knowledge. Finally, we explore the generalization of the discovered mechanisms across various syllogistic schemes and model sizes, finding that the identified circuit is sufficient and necessary for all the schemes on which the model achieves high downstream accuracy ($\geq$ 60%). Overall, our findings suggest that LMs indeed learn transferable content-independent reasoning mechanisms, but that, at the same time, such mechanisms do not involve generalisable and abstract logical primitives, being susceptible to contamination by the same world knowledge acquired during pre-training.
Introduction
Transformer-based Language Models (LMs) have led to remarkable results across various natural language processing tasks (Radford et al. 2018, 2019; Brown et al. 2020; OpenAI et al. 2024; Jason et al. 2022; Bubeck et al. 2023). This success has catalyzed research interest in systematically exploring the reasoning capabilities emerging during pre-training (Clark, Tafjord, and Richardson 2020). Recent findings suggest that logical reasoning abilities may emerge in large-scale models (Rae et al. 2022; Kojima et al. 2022; Wei et al. 2022) or through transfer learning on specialized datasets (Betz, Voigt, and Richardson 2021). However, ongoing debates question whether these models apply systematic inference rules or reuse superficial patterns learned during pre-training (Talmor et al. 2020; Kassner, Krojer, and Schütze 2020; Wu et al. 2024). This controversy underscores the need for a deeper understanding of the low-level logical inference mechanisms in LMs (Rozanova, Valentino, and Freitas 2024; Rozanova et al. 2023, 2022).
@@ Introduction
To improve our understanding of the internal mechanisms, this paper focuses on Mechanistic Interpretability (Olah et al. 2020; Nanda et al. 2023), aiming to discover the core circuit responsible for content-independent syllogistic reasoning. In particular, we focus on categorical syllogisms with universal affirmative quantifiers (i.e., AAA-1, Barbara) motivated by two key factors. First, as observed in natural logic studies (MacCartney and Manning 2007), this form of syllogistic reasoning is prevalent in everyday language. Therefore, it is likely that LMs are exposed to such reasoning schema during pre-training. Second, AAA-1 is a form of unconditionally valid syllogism independent of the premises’ truth condition (Holyoak and Morrison 2005). This characteristic offers a deterministic and scalable task design, other than allowing us to investigate the disentanglement between reasoning and knowledge representation.
Through the adoption of Activation Patching (Meng et al. 2022) and embedding space analysis (i.e., Logit Lens) (Nostalgebraist 2020; Geva et al. 2022; Dar et al. 2023), we investigate the following main research questions: RQ1: What is the content-independent syllogistic reasoning mechanism internalized in LMs during pre-training?; RQ2: Are content-independent mechanisms disentangled from specific world knowledge and belief biases?; RQ3: Does the core reasoning mechanism generalize across model sizes and different syllogistic schemes?
@@
  1. 1.
    The circuit analysis reveals that auto-regressive LMs develop specific inference mechanisms during pre-training, finding evidence supporting a three-stage mechanism for syllogistic reasoning: (1) naive recitation of the first premise; (2) suppression of duplicated middle-term information; and (3) mediation towards the correct output through the interplay of attention heads.
  2. 2.
    Further experiments on circuit transferability demonstrate that the identified mechanism is still necessary for reasoning on syllogistic schemes instantiated with commonsense knowledge. However, a deeper analysis suggests that specific belief biases acquired during pre-training might play an important role, contaminating the content-independent circuit mechanism with additional attention heads responsible for encoding general and contextualized world knowledge.
  3. 3.
    We found that the identified circuit is sufficient and necessary for all the syllogistic schemes in which the model achieves high downstream accuracy ($\geq$ 60%). This result suggests that LMs learn content-independent reasoning mechanisms transferable across abstract schemes.
  4. 4.
    The intervention results on models with different sizes show similar suppression mechanism patterns and information flow. However, we found evidence that the contribution of attention heads becomes more complex with model sizes, further supporting the hypothesis of increasing contamination from external world knowledge.
Our following study is conducted using TransformerLens (Nanda and Bloom 2022) on an Nvidia A100 GPU with 80GB of memory. The dataset and code to reproduce our experiments are available online to encourage future work in the field [^1] http://anonynous˙url.com.
Methodology
Our main research objective is to discover and interpret the core mechanisms adopted by auto-regressive Language Models (LMs) when performing content-independent syllogistic reasoning. To this end, we present a methodology that consists of 3 main stages. First, we define a syllogistic completion task that can be instrumental for our analysis. Second, we leverage the syllogistic completion task to implement a circuit discovery pipeline on a syllogistic schema instantiated only with symbolic variables (Table 1, Symbolic). Third, we investigate the generalization of the identified circuit on concrete schemes instantiated with commonsense knowledge and explore how the internal behavior varies with different schemes and model sizes.
@@
Table 1: Examples of the same syllogistic schema (i.e., AAA-1, Barbara) instantiated with different content (with the expected completion in bold). The ability to complete the syllogism is a function of the reasoning schema, independent of the specific instantiation or truth condition of the premises.
Figure 2: The conceptual pipeline of the symbolic circuit analysis on the syllogism completion task. Figures (a)–(e) illustrate the stages involved in processing a syllogistic argument: argument formalization (a), construction of clean (b) and corrupted inputs (c), activation patching (d) and embedding space analysis (i.e., Logit Lens) (e). The uncovered symbolic circuit is subsequently evaluated to determine whether the core reasoning schema operates independently from world knowledge or belief bias.
Syllogism Completion Task
We design a syllogism completion task for assessing the reasoning abilities of LMs, building upon established approaches in the literature (Betz, Voigt, and Richardson 2021; Wu et al. 2023). In particular, we focus on categorical syllogisms with universal affirmative quantifiers (i.e., AAA-1, Barbara) because of their frequency in natural language and their content-independent reasoning property.
The syllogistic argument is typically composed of two premises ($\mathcal{P}_{1},\mathcal{P}_{2}$) and a conclusion ($\mathcal{C}$), with $s$ and $p$ denoting the subject and predicate terms in the conclusion (e.g. men and mortal), and $m_{1}$ and $m_{2}$ (e.g., humans) denoting the middle terms in the two premises, with $m_{1}\equiv m_{2}$ in the case of the AAA-1 syllogism.
To evaluate syllogistic reasoning in LMs, we formalize the completion task as masking language modeling, removing the final token $p$ from the conclusion (e.g., All men are humans. All humans are mortal. Therefore, all men are), and comparing the probability assigned by the LM to $p$ (e.g., mortal) and the middle term $m$ (e.g., humans). In general, an LM is successful in the completion of a task if the following condition applies:
@@
We conduct experiments using two distinct datasets, symbolic and non-symbolic, to derive comparative implications for model reasoning capabilities. The symbolic dataset is constructed by randomly sampling triples (e.g., A, B, and C) from the set of 26 uppercase letters of the English alphabet. On the other side, the non-symbolic dataset is constructed by replacing the letters with words while preserving the syllogistic schema. Here, we generate two different non-symbolic sets: a belief-consistent set containing true premises and a belief-inconsistent set containing false premises (see Table 1). To guarantee the truth condition of the premises, we leverage GenericsKB (Bhakthavatsalam, Anastasiades, and Clark 2020), a knowledge base containing statements about commonsense knowledge. The detailed generation process is in Appendix B.
Circuit Discovery
Our main objective is to find a circuit $\mathcal{C}$ for syllogistic reasoning in a Language Model $\mathcal{M}$. A circuit $\mathcal{C}$ can be defined as a subset of the original model $\mathcal{M}$ that is both sufficient and necessary for achieving the original model performance on the syllogistic completion task. In order to identify a circuit, we employ Activation Patching together with circuit ablation methods (Meng et al. 2022; Vig et al. 2020).
@@
  1. (1)
    Clean Run: For the target input $(s,m_{1},m_{2},p)$, we record the baseline logit difference, $\delta_{+}(p,m)$ from the forward pass output of the model (Figure 2(1)).
  2. (2)
    Corrupted Run: We re-run the model on a corrupted input after applying a specific intervention (e.g., changing the middle term $m_{2}$) and record the logit difference $\delta_{-}(p,m)$ (Figure 2(3)).
  3. (3)
    Intervened Run: We run the model with the corrupted inputs again and replace the corrupted activations with those from the clean runs to compute the response from the remaining components and measure the causal impact of the intervention. Here, we measure the adjusted logit difference $\delta_{p}(p,m)$ (Figure 2(2)).
  4. (4)
    Quantification: We quantify the causal impact of each intervention using a patching score $\mathcal{S}$ following (Heimersheim and Nanda 2024) as shown in Figure 2(4). This score is further normalized to $[-1,1]$.
We complement Activation Patching with known methods for analysing hidden activations in transformers, including Logit Lens (Nostalgebraist 2020) with an input-agnostic approach (Dar et al. 2023; Hanna, Liu, and Variengien 2024). Additional technical details can be found in Appendix C.
Interventions.
@@
  1. (1)
    Middle-Term Corruption: To investigate the reasoning mechanism employed in the syllogism completion task, we primarily focus on the transitive property of the middle term, $(m_{1}\rightarrow m_{2})\wedge(m_{1}\equiv m_{2})$. We hypothesize that disrupting the transitive property will localize the component responsible for syllogistic reasoning. Therefore, our intervention method replaces the second middle term $m_{2}$ with an unseen symbol $m_{2}^{\prime}$, breaking the equality and effectively corrupting the validity of the reasoning, pivoting the correct answer towards $m$: $(s,m_{1},m_{2},p)\rightarrow(s,m_{1},m_{2}^{\prime},p)$.
  2. (2)
    All-Term Corruption: We examine how term-related information flows to the last position for the final prediction to identify potential mover heads. To this end, we replace the original terms $(s,m_{1},m_{2},p)$ with different terms $(s^{\prime},m_{1}^{\prime},m_{2}^{\prime},p^{\prime})$ while keeping the completion answers unchanged. We hypothesize that if the mover heads carry the $p$ related information over $m$ related information for a valid prediction, the logit difference will increase; otherwise, it will decrease.
Figure 3: Accuracy on the syllogism completion task across all scales of GPT-2. We observe a phase change from small to medium models in overall performance.
Circuit Ablation.
@@
Empirical Evaluation
Figure 4: Comprehensive results of the symbolic circuit analysis. (a) Attention output patching results and (b) residual stream patching results in the middle-term intervention setup. (c) OV circuit logit lens results for $\mathbf{h}_{11.10}$, with input and output comprising 26 uppercase letters. (d) Residual stream patching results in the all-term corruption setup. (e)-(h) Average attention patching weights across the batch: (e) weights for token position $[p]$ attending, and (f)-(h) weights for the last token position attending. For clarity, the y-axis legend denotes tokens appearing between terms, separated by a ‘-’ symbol, with averaged values.
Model Selection.
We select the LM for the circuit analysis based on a trade-off between model size and performance. To this end, we measure the accuracy of GPT-2 (Radford et al. 2019) with different sizes (117M, 345M, 762M, and 1.5B) on the syllogism completion task. Specifically, we aim to identify potential phase transition points where significant performance changes emerge (Jason et al. 2022; Kaplan et al. 2020). As shown in Figure 3, we observe a notable phase transition from small to medium sizes, where average performance increases by 139.97% in the symbolic setting and 119.21% in the non-symbolic setting. Given these results, we select GPT-2 Medium for circuit discovery and subsequently test the compatibility with different model sizes (See Appendix E for the model architecture). For all experiments, we use 90 samples each for symbolic and non-symbolic arguments to balance the depth of analysis with computational efficiency.
Localization of Transitive Reasoning Mechanisms
@@
Figure 4(a-b) presents the results of the middle-term intervention targeting attention heads and residual stream. Figure 4(a), in particular, reveals the positive role of heads $\mathbf{h}_{11.10}$ and $\mathbf{h}_{19.1}$ (where $\mathbf{h}_{l.k}$ denotes the $k$th head in layer $l$). Moreover, we observe an information propagation pattern from the $[m_{2}]$ position to the $[p]$ position (Figure 4(b)). These observations allow us to hypothesize that information from $[m_{2}]$ is conveyed to $[p]$ on the residual stream by attention head $\mathbf{h}_{11.10}$, which exhibits a strong patching score around the layer responsible for the propagation. To verify this hypothesis, we further investigate the attention weights between $[p]$ and $[m_{2}]$. As expected, $[m_{2}]$ is the position most highly attended by $[p]$, with an average attention weight of $0.15\pm 0.07$ (Figure 4(e)). These results suggest that information is indeed moved from $[m_{2}]$ to $[p]$ on the residual stream subspace, with $\mathbf{h}_{11.10}$ playing a crucial role in this mechanism.
@@
We further investigate how information from $[m_{2}]$ is moved to $[p]$, positing that this relates to the model’s internal reasoning mechanism. We notice that at position $[p]$, the model can observe the complete AAA-1 syllogistic structure (Figure 2(a)) with middle-term duplication. We hypothesize that this structural information for reasoning is collected at one position, given that information refinement occurs at the specific position such as last token (Hanna, Liu, and Variengien 2024; Stolfo, Belinkov, and Sachan 2023). To verify this, we trace the information flow of head $\mathbf{h}_{11.10}$ using path patching (Wang et al. 2023). Our results show that $\mathbf{h}_{11.10}$ operates based on several induction heads (Elhage et al. 2021) ($\mathbf{h}_{5.8}$, $\mathbf{h}_{6.1}$, $\mathbf{h}_{6.15}$ and $\mathbf{h}_{7.2}$) formed at $[m_{2}]$. These heads attend to the $[[m_{1}]+1]$ token due to the $m_{1}\equiv m_{2}$ conditioned matching operation, likely containing $m_{1}$-related information from previous token heads. We conclude, therefore, that $m_{1}$ and $m_{2}$ information are aggregated at position $[p]$. Additional details of the path patching results are available in the Appendix F.
@@
To better understand the internal mechanism occurring at the $[p]$ position, we investigate attention head $\mathbf{h}_{11.10}$. Having previously examined the attention pattern (composed of attention weights), we now focus on the attention value and output by analyzing the OV circuit ($W_{V}W_{O}$(Elhage et al. 2021) using the logit lens method (Nostalgebraist 2020). Interestingly, the result in Figure 4(c) shows a clear negative diagonal pattern suggesting that $\mathbf{h}_{11.10}$ strongly suppresses the logit when attending to the same token as the corresponding output. Given our previous findings that $\mathbf{h}_{11.10}$ reads information from the subspace of token $[m_{2}]$, we conclude that it applies a suppressive mechanism to $m$-related information and writes it back to the residual stream’s subspace at $[p]$. If later heads carry this information to the last token, this mechanism becomes crucial for the model to arrive at the correct answer. For simplicity, we name head $\mathbf{h}_{11.10}$ as $m$-suppression head.
Localization of Term-Related Information Flow

### v2

@@ Abstract
Recent studies on logical reasoning in Language Models (LMs) have sparked a debate on whether they can learn systematic reasoning principles during pre-training or merely exploit superficial patterns in the training data. This paper presents a mechanistic interpretation of syllogistic reasoning in LMs to advance the understanding of internal dynamics. Specifically, we present a methodology for circuit discovery aimed at interpreting content-independent reasoning mechanisms. Through two distinct intervention methods, we uncover a sufficient and necessary circuit involving middle-term suppression that elucidates how LMs transfer information to derive valid conclusions from premises. Furthermore, we investigate how belief biases manifest in syllogistic reasoning, finding evidence of partial contamination from additional attention heads responsible for encoding commonsense and contextualized knowledge. Finally, we explore the generalization of the discovered mechanisms across various syllogistic schemes, model sizes and architectures, finding that the identified circuit is sufficient and necessary for the schemes on which the models achieve high downstream accuracy ($\geq$ 60%), and that the activation patterns apply to models of different families. Overall, our findings suggest that LMs indeed learn transferable content-independent reasoning mechanisms, but that, at the same time, such mechanisms do not involve generalizable and abstract logical primitives, being susceptible to contamination by the same world knowledge acquired during pre-training.
A Mechanistic Interpretation of Syllogistic Reasoning in
Auto-Regressive Language Models
Geonhee Kim1, Marco Valentino3, André Freitas1,2,3 1Department of Computer Science, University of Manchester, UK 2National Biomarker Centre (NBC), CRUK Manchester Institute, UK 3 Idiap Research Institute, Switzerland
1 Introduction
@@
Transformer-based Language Models (LMs) have led to remarkable results across various natural language processing tasks (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI et al., 2024; Jason et al., 2022; Bubeck et al., 2023). This success has catalyzed research interest in systematically exploring the reasoning capabilities emerging during pre-training (Clark et al., 2020). Recent findings suggest that logical reasoning abilities may emerge in large-scale models (Rae et al., 2022; Kojima et al., 2022; Wei et al., 2022) or through transfer learning on specialized datasets (Betz et al., 2021). However, ongoing debates question whether these models apply systematic inference rules or reuse superficial patterns learned during pre-training (Talmor et al., 2020; Kassner et al., 2020; Wu et al., 2024). This controversy underscores the need for a deeper understanding of the low-level logical inference mechanisms in LMs (Rozanova et al., 2024, 2023, 2022).
To improve our understanding of the internal mechanisms, this paper focuses on Mechanistic Interpretability (Olah et al., 2020; Nanda et al., 2023), aiming to discover the core circuit responsible for syllogistic reasoning. In particular, we focus on categorical syllogisms with universal affirmative quantifiers (i.e., AAA-1, Barbara) motivated by two key factors. First, as observed in natural logic studies (MacCartney and Manning, 2007), this form of syllogistic reasoning is prevalent in everyday language. Therefore, it is likely that LMs are exposed to such reasoning schema during pre-training. Second, AAA-1 is a form of unconditionally valid syllogism independent of the premises’ truth condition (Holyoak and Morrison, 2005). This characteristic offers a deterministic and scalable task design, other than allowing us to investigate the disentanglement between reasoning and knowledge representation Wysocka et al. (2024); Bertolazzi et al. (2024); Lampinen et al. (2024).
Through mechanistic interpretability techniques such as Activation Patching (Meng et al., 2022) and embedding space analysis (i.e., Logit Lens) (Nostalgebraist, 2020; Geva et al., 2022; Dar et al., 2023), we investigate the following main research questions: RQ1: How is the content-independent syllogistic reasoning mechanism internalized in LMs during pre-training?; RQ2: Are content-independent mechanisms disentangled from specific world knowledge and belief biases?; RQ3: Does the core reasoning mechanism generalize across model sizes and syllogistic schemes?
@@
1. The circuit analysis reveals that LMs develop specific inference mechanisms during pre-training, finding evidence supporting a three-stage mechanism for syllogistic reasoning: (1) naive recitation of the first premise; (2) suppression of duplicated middle-term information; and (3) mediation towards the correct output through the interplay of attention heads.
2. Further experiments on circuit transferability demonstrate that the identified mechanism is still necessary for reasoning on syllogistic schemes instantiated with commonsense knowledge. However, a deeper analysis suggests that specific belief biases acquired during pre-training might contaminate the content-independent circuit mechanism with additional attention heads responsible for encoding contextualized world knowledge.
3. We found that the identified circuit is sufficient and necessary for all the syllogistic schemes in which the model achieves high downstream accuracy ($\geq$ 60%). This result suggests that LMs learn reasoning mechanisms that are transferable across different schemes.
4. The intervention results on models with different architectures and sizes (i.e., GPT-2 (Radford et al., 2019), Pythia (Biderman et al., 2023), Llama (Grattafiori et al., 2024), Qwen (Yang et al., 2024)) show similar suppression mechanism patterns and information flow. However, we found evidence that the contribution of attention heads becomes more complex with model sizes, further supporting the hypothesis of increasing contamination from external world knowledge.
@@
Table 1: Examples of a syllogistic schema (i.e., AAA-1, Barbara). The logical validity of a conclusion is only a function of the reasoning schema, being independent of the specific variables or truth condition of the premises.
Figure 2: The conceptual pipeline of the circuit discovery methodology. Figures (a)–(e) illustrate the stages in processing a syllogistic argument: formalization (a), construction of clean (b) and corrupted inputs (c), activation patching (d) and embedding space analysis (i.e., Logit Lens) (e). The uncovered symbolic circuit is evaluated to determine whether the core reasoning schema operates independently from world knowledge or belief bias.
2 Methodology
Our main research objective is to discover and interpret the core mechanisms adopted by auto-regressive Language Models (LMs) when performing content-independent syllogistic reasoning. To this end, we present a methodology that consists of 3 main stages. First, we define a syllogistic completion task that can be instrumental for our analysis. Second, we leverage the syllogistic completion task to implement a circuit discovery pipeline on a syllogistic schema instantiated only with symbolic variables (Table 1, Symbolic). Third, we investigate the generalization of the identified circuit on concrete schemes instantiated with commonsense knowledge and explore how the internal behavior varies with different schemes and models.
2.1 Syllogism Completion Task
We design a syllogism completion task for assessing the reasoning abilities of LMs, building upon established approaches in the literature (Betz et al., 2021; Wu et al., 2023). In particular, we focus on categorical syllogisms with universal affirmative quantifiers (i.e., AAA-1, Barbara) because of their frequency in natural language and their content-independent reasoning property. The syllogistic argument is typically composed of two premises ($\mathcal{P}_{1},\mathcal{P}_{2}$) and a conclusion ($\mathcal{C}$), with $s$ and $p$ denoting the subject and predicate terms in the conclusion (e.g. men and mortal), and $m_{1}$ and $m_{2}$ (e.g., humans) denoting the middle terms in the two premises, with $m_{1}\equiv m_{2}$ in the case of the AAA-1 syllogism.
To evaluate syllogistic reasoning in LMs, we formalize the completion task as language modeling, removing the final token $p$ from the conclusion (e.g., All men are humans. All humans are mortal. Therefore, all men are), and comparing the probability assigned by the LM to $p$ (e.g., mortal) and the middle term $m$ (e.g., humans). In general, an LM is successful in the completion of a task if the following condition applies:
@@ 2.1 Syllogism Completion Task
We conduct experiments using two distinct datasets, symbolic and non-symbolic, to derive comparative implications for model reasoning capabilities. The symbolic dataset is constructed by randomly sampling triples (e.g., A, B, and C) from the set of 26 uppercase letters of the English alphabet. On the other side, the non-symbolic dataset is constructed by replacing the letters with words while preserving the syllogistic schema. Here, we generate two different non-symbolic sets: a belief-consistent set containing true premises and a belief-inconsistent set containing false premises to address RQ2 (see Table 1). To guarantee the truth condition of the premises, we leverage GenericsKB (Bhakthavatsalam et al., 2020), a knowledge base containing statements about commonsense knowledge. The detailed generation process is in Appendix B.
2.2 Circuit Discovery
Our main objective is to find a circuit $\mathcal{C}$ for syllogistic reasoning in a Language Model $\mathcal{M}$. A circuit $\mathcal{C}$ can be defined as a subset of the original model $\mathcal{M}$ that is both sufficient and necessary for achieving the original model performance on the syllogistic completion task. In order to identify a circuit, we employ Activation Patching together with circuit ablation (Meng et al., 2022; Vig et al., 2020).
@@ 2.2 Circuit Discovery
(1) Clean Run.
For the target input $(s,m_{1},m_{2},p)$, we record the baseline logit difference, $\delta_{+}(p,m)$ from the forward pass output of the model (Figure 2(1)).
(2) Corrupted Run.
We re-run the model on a corrupted input after applying a specific intervention (e.g., changing the middle term $m_{2}$) and record the logit difference $\delta_{-}(p,m)$ (Figure 2(3)).
(3) Intervened Run.
We run the model with the corrupted inputs again and replace the corrupted activations with those from the clean runs to compute the response from the remaining components and measure the causal impact of the intervention. Here, we measure the adjusted logit difference $\delta_{p}(p,m)$ (Figure 2(2)).
(4) Quantification.
We quantify the causal impact of each intervention using a patching score $\mathcal{S}$ following (Heimersheim and Nanda, 2024) as shown in Figure 2(4). This score is further normalized to $[-1,1]$.
We complement Activation Patching with known methods for analysing hidden activations in transformers, including Logit Lens (Nostalgebraist, 2020) with an input-agnostic approach (Dar et al., 2023; Hanna et al., 2024). Additional technical details can be found in Appendix C.
Figure 3: Results of the symbolic circuit analysis. (a) Attention output patching and (b) residual stream patching obtained via middle-term intervention. (c) OV circuit logit lens results for attention head $\mathbf{h}_{11.10}$, with input and output comprising 26 uppercase letters. (d) Residual stream patching obtained via all-term corruption. (e)-(h) Average attention weights across the batch: (e) weights for token position $[p]$, and (f)-(h) weights for the last token position. For clarity, the symbol ‘-’ on the y-axis denotes the averaged values for tokens appearing between terms.
2.3 Interventions
@@
(1) Middle-Term Corruption.
To investigate the reasoning mechanism employed in the syllogism completion task, we primarily focus on the transitive property of the middle term, $(m_{1}\rightarrow m_{2})\wedge(m_{1}\equiv m_{2})$. We hypothesize that disrupting the transitive property will localize the component responsible for syllogistic reasoning. Therefore, our intervention method replaces the second middle term $m_{2}$ with an unseen symbol $m_{2}^{\prime}$, breaking the equality and effectively corrupting the validity of the reasoning, pivoting the correct answer towards $m$: $(s,m_{1},m_{2},p)\rightarrow(s,m_{1},m_{2}^{\prime},p)$.
(2) All-Term Corruption.
We examine how term-related information flows to the last position for the final prediction to identify potential mover heads. To this end, we replace the original terms $(s,m_{1},m_{2},p)$ with different terms $(s^{\prime},m_{1}^{\prime},m_{2}^{\prime},p^{\prime})$ while keeping the answers unchanged. We hypothesize that if heads carry the information preferring $p$ over $m$ for the valid prediction, the logit difference will increase; otherwise, it will decrease.
2.4 Circuit Ablation.
@@
3 Empirical Evaluation
Empirical Setup.
We select the LM for the circuit analysis based on a trade-off between model size and performance. To this end, we measure the accuracy of GPT-2 (Radford et al., 2019) with different sizes (117M, 345M, 762M, and 1.5B) on the syllogism completion task. Specifically, we aim to identify potential phase transition points where significant performance changes emerge (Jason et al., 2022; Kaplan et al., 2020). As shown in Figure 6(a), we observe a notable transition from small to medium sizes, where averaged performance in three different datasets increases by 439.06%. The shift is more apparent under the logit difference metric (See Section 3.8 and Figure 6(b)). Given these results, we select GPT-2 Medium for circuit discovery and subsequently test the compatibility with different model sizes and architectures (Section 3.7 and 3.8).
3.1 Transitive Reasoning Mechanisms
@@
Figure 3(a-b) presents the results of the middle-term intervention targeting attention heads and residual stream. Figure 3(a), in particular, reveals the positive role of heads $\mathbf{h}_{11.10}$ and $\mathbf{h}_{19.1}$ (where $\mathbf{h}_{l.k}$ denotes the $k$th head in layer $l$). Moreover, we observe an information propagation pattern from the $[m_{2}]$ position to the $[p]$ position (Figure 3(b)). These observations allow us to hypothesize that information from $[m_{2}]$ is conveyed to $[p]$ on the residual stream by attention head $\mathbf{h}_{11.10}$, which exhibits a strong patching score around the layer responsible for the propagation. To verify this hypothesis, we further investigate the attention weights between $[p]$ and $[m_{2}]$. As expected, $[m_{2}]$ is the position most highly attended by $[p]$, with an average attention weights of $0.15\pm 0.07$ (Figure 3(e)). These results suggest that information is indeed moved from $[m_{2}]$ to $[p]$ on the residual stream subspace, with $\mathbf{h}_{11.10}$ playing a crucial role in this mechanism.
@@
We further investigate how information from $[m_{2}]$ is moved to $[p]$, positing that this relates to the model’s internal reasoning mechanism. We notice that at position $[p]$, the model can observe the complete AAA-1 syllogistic structure (Figure 2(a)) with middle-term duplication. We hypothesize that this structural information for reasoning is collected at one position, given that information refinement occurs at the specific position such as last token (Hanna et al., 2024; Stolfo et al., 2023). To verify this, we trace the information flow of head $\mathbf{h}_{11.10}$ using path patching (Wang et al., 2023). Our results show that $\mathbf{h}_{11.10}$ operates based on several induction heads (Elhage et al., 2021) ($\mathbf{h}_{5.8}$, $\mathbf{h}_{6.1}$, $\mathbf{h}_{6.15}$ and $\mathbf{h}_{7.2}$) formed at $[m_{2}]$. These heads attend to the $[[m_{1}]+1]$ token due to the $m_{1}\equiv m_{2}$ conditioned matching operation, likely containing $m_{1}$-related information from previous token heads. We conclude, therefore, that $m_{1}$ and $m_{2}$ information are aggregated at position $[p]$. Additional details of the path patching results are available in the Appendix G.
@@
To better understand the internal mechanism occurring at the $[p]$ position, we investigate attention head $\mathbf{h}_{11.10}$. Having previously examined the attention pattern (composed of attention weights), we now focus on the attention value and output by analyzing the OV circuit ($W_{V}W_{O}$(Elhage et al., 2021) using the logit lens method (Nostalgebraist, 2020). Interestingly, the result in Figure 3(c) shows a clear negative diagonal pattern suggesting that $\mathbf{h}_{11.10}$ strongly suppresses the logit when attending to the same token as the corresponding output. Given our previous findings that $\mathbf{h}_{11.10}$ reads information from the subspace of token $[m_{2}]$, we conclude that it applies a suppressive mechanism to $m$-related information and writes it back to the residual stream’s subspace at $[p]$. If later heads carry this information to the last token, this mechanism becomes crucial for the model to arrive at the correct answer. For simplicity, we name head $\mathbf{h}_{11.10}$ as $m$-suppression head.
3.2 Term-Related Information Flow